# Phase diagram of the hard-sphere potential model in three and four dimensions using a pseudo-hard-sphere potential

### v1

@@ Abstract
The hard-sphere potential has become a cornerstone in the study of both molecular and complex fluids. Despite its mathematical simplicity, its implementation in fixed time-step molecular simulations remains a formidable challenge due to the discontinuity at contact. To circumvent the issues associated with the ill-defined force at contact, a continuous potential—referred to here as the pseudo-hard-sphere (pHS) potential—has recently been proposed [J. Chen, Phys. 149, 164907 (2018)]. This potential is constructed to match the second virial coefficient of the hard-sphere potential and is expected to mimic its thermodynamic properties. However, this hypothesis has only been partially validated within the fluid region of the phase diagram for hard-sphere dispersions in two and three dimensions. In this contribution, we examine the ability of the continuous pHS potential to reproduce the equation of state of a hard-sphere fluid, not only in the fluid phase but also across the fluid-solid coexistence region. Our focus is primarily on hard-sphere systems in three and four dimensions. We compare the results obtained from Brownian dynamics simulations of the pHS potential with those derived from refined event-driven simulations of the corresponding hard-sphere potential. Furthermore, we provide a comparative analysis with theoretical equations of state based on both mean-field and integral equation approximations.
@@ I Introduction
The hard-sphere (HS) model remains the reference system for liquids and soft matter due to its simplicity, physical relevance, and rich behavior. It has been extensively studied over the years [1, 2, 3] through computer simulations [4, 5] and is commonly employed as a reference system in perturbation-based thermodynamic approximations [6, 7]. The continued interest in the HS model stems from the extensive knowledge available and the feasibility of experimental realizations [8, 9, 10, 11], made possible by recent advances in microscopy and scattering techniques. These developments, together with complementary computer simulations [12] and theoretical studies using integral equation theory [13], establish the HS model as a robust framework for testing theories across the full spectrum of current soft matter physics research.
@@
Computer simulations are among the principal techniques in soft matter research and have been extensively employed to study the phase diagram of the HS model [14, 12, 3]. However, the inherent discontinuity of the HS model renders continuous time integration methods unsuitable [15, 5]. In particular, Brownian dynamics simulations [16, 2] cannot be directly applied to the HS model, necessitating the use of specialized event-driven algorithms [17]. Recently, Jover et al. [18] proposed mapping the HS model to a continuous interaction potential. In their approach, the compressibility factor was compared against the well-established Carnahan–Starling (CS) equation of state [2], and the parameters of the potential were fixed when the compressibility from simulations matched the analytical CS expression. Rather than relying on trial and error, the work of Báez et al. [19] employed the extended law of corresponding states [20] to determine a set of potential parameters by directly matching the second virial coefficient of the HS potential with that of the continuous interaction potential. This procedure yields a pseudo hard-sphere (pHS) interaction potential with parameters that are independent of thermodynamic conditions, such as density.
@@
The mapping of the HS model to a continuous potential has been demonstrated to reproduce the thermodynamics of the real HS one-component fluid [18], as well as those of binary and polydisperse mixtures [21]. The pHS model has recently been employed to inverse design the self-assembly of exotic crystalline structures [22]. In addition, computer simulations have been used to investigate the dynamics and transport phenomena of the pHS model [23, 24], and the interaction potential has found applications in modeling active matter [25] and in evaluating depletion forces in binary and ternary colloidal mixtures [26, 27]. These examples underscore the robustness of the model and its wide range of applications in systems where short-ranged, hard-like interactions are required. Furthermore, the pHS interaction potential has been successfully applied within the Ornstein-Zernike framework to reproduce thermodynamic properties, particularly in studies of polydisperse fluids [21].
@@
The aim of this work is to demonstrate the applicability of the pHS model across the full range of the three-dimensional fluid and solid phases, and to determine the coexistence densities at the fluid–solid transition using Brownian dynamics simulations. We solve the Ornstein–Zernike equation with various closure relations to obtain the liquid equation of state over a range of densities. Furthermore, we test the hypothesis proposed by Báez et al. [19] that the potential is applicable in arbitrary spatial dimensions by performing computer simulations and solving the Ornstein–Zernike equation in four dimensions. Our results show excellent agreement with mean-field equations of state, previously reported simulation data used for fitting semi-phenomenological equations of state, and solutions of the Ornstein–Zernike equation.
@@
Following this Introduction, the manuscript is organized as follows. In Sec.II, we introduce the pHS model, outline its derivation, and specify the parameters used for the interaction potential throughout this work. Sec.III provides a detailed description of the methods employed for the computer simulations and the numerical solution of the Ornstein–Zernike equation, including the relevant parameters. In Sec.V, we present and discuss the computer simulation results for the three- and four-dimensional systems. Sec.IV is dedicated to the presentation and discussion of the results obtained from the Ornstein–Zernike equation. Finally, Sec. VI concludes the work with an overview of the main findings and perspectives based on the results presented.
@@ II Pseudo-hard sphere potential revisited
The work of Jover et al. [18] introduced a continuous interaction potential, which is a generalized Weeks-Chandlers-Andersen (WCA), to reproduce the CS equation of state. The WCA interaction potential is defined as [18]
$u_{WCA}(r)=\left\{\begin{array}[]{cc}A\varepsilon\left[\left(\frac{\sigma}{r}\right)^{\lambda}-\left(\frac{\sigma}{r}\right)^{\lambda-1}\right]+\varepsilon&r<\sigma B,\\ 0&r\geq\sigma B,\end{array}\right.$ (2)
@@
$A=\lambda\quantity(\frac{\lambda}{\lambda-1})^{\lambda-1}\quad,\quad B=\quantity(\frac{\lambda}{\lambda-1})$ (3)
The exponent $\lambda$ is related to the stiffness of the WCA potential, and $\epsilon$ is the energy parameter that measures the repulsion strength between a pair of particle. Typically, $\epsilon$ is used to define the reduced temperature $k_{B}T/\epsilon$, where $k_{B}$ is the Boltzmann constant and $T$ the absolute temperature.
Later, in the work of Báez et al. [19], it was shown that following the so-called extended law of corresponding states [20], the assumption that the second virial coefficient of the continuous potential $B_{2}^{\mathrm{WCA}}$ must be equal to the second virial coefficient of the HS interaction potential, $B_{2}^{\mathrm{HS}}$, will yield a set of values for $\epsilon/k_{B}T$ in terms of $\lambda$ that successfully maps the HS model to a continuous one in any spatial dimension, $d$. In principle, one can choose an arbitrary value for $\lambda$, however, we follow the work of Báez et al. [19] and set the value of $\lambda=50$. For each spatial dimension, a value of the reduced temperature will be obtained. In particular, for three dimensions we use $k_{B}T_{\mathrm{3D}}/\epsilon=1.4737$, and for four dimensions we use $k_{B}T_{\mathrm{4D}}/\epsilon=1.4803$  [19].
@@
The equation of state in $d$ dimensions is related to the contact value of the pair correlation function $g(\sigma)$ by [28]
$Z=\frac{\beta P}{\rho}=1+B_{2}\rho g(\sigma)\,,$ (5)
@@
III Integral equations theory and computer simulation details
@@ III.1.1 Three dimensional simulations
For the three dimensional system of hard spheres, we simulate a system of $N=20376$ particles in a fixed volume $V$, and energy $E$. We perform event-driven molecular dynamics (EDMD) using the algorithm of  Smallenburg [30]. Initial configurations are obtained by starting in a dilute state at the desired density, and then performing an EDMD simulation in which the particle diameters grow until the desired packing fraction is reached. After the packing fraction was reached, the system is equilibrated for at least $10^{7}\tau_{\mathrm{MD}}$ and data collection for the pressure was done for another $10^{8}\tau_{\mathrm{MD}}$, with $\tau_{\mathrm{MD}}=\sigma\sqrt{m/k_{B}T}$ the simulation time unit, $m$ the mass of the particle, which is the same for all particles. Pressure is measured using the virial expression [30]. For the solid branch, all details remain the same except for the initial configuration, for which a face centered cubic (FCC) lattice was used for each of the target packing fractions. We estimate the average value of the pressure and the error of the measurement using block analysis [31].
@@
The three dimensional pHS model was simulated using the HOOMD-blue code [32]. Brownian dynamics (BD) simulations were performed using the standard Euler scheme to solve the equations of motion for the particles in the canonical ($NVT$) ensemble [16, 33]. For the fluid branch of the equation of state, we simulate a total of $N=32000$ particles at a fixed reduced temperature of $k_{B}T/\epsilon=1.4737$ following the result from Báez et al. [19]. We use a time step of $\Delta\tau_{\mathrm{BD}}=10^{-5}$, where $\tau_{\mathrm{BD}}=\sigma^{2}/D_{0}$ is the Brownian time unit, and $D_{0}=k_{B}T/3\pi\eta_{0}\sigma$ is the free-particle diffusion coefficient, with $\eta_{0}$ being the zero-frequency shear viscosity. We initialize the system at a low density and slowly compress the simulation box over a total simulation time of $10^{5}\tau_{\mathrm{BD}}$ until the target packing fraction has been reached. Once the desired density has been reached, the system is equilibrated for at least $5\times 10^{7}\tau_{\mathrm{BD}}$, and an additional production run of $10^{7}\tau_{\mathrm{BD}}$ is used to collect data for the pressure, which is computed through the virial expression for a pairwise interaction potential [5]. For the solid branch of the equation of state of the pHS model, we use a smaller time step of $\Delta\tau_{\mathrm{BD}}=5\times 10^{-7}$, and initialize the system in the FCC configuration for the desired target packing fraction.
@@
To determine the fluid-crystal coexistence properties in the three dimensional pHS model, we calculate the coexistence pressure using the $NVT$ ensemble method introduced by Smallenburg et al. [34]. Below, we provide a brief summary of the method, but we direct readers to the original publication for a complete description of the method. The coexistence pressure in a fluid-crystal system is estimated by performing direct coexistence simulations in the $NVT$ ensemble to evaluate the system’s pressure tensor. The pressures of the fluid and solid are then compared, as true coexistence requires the pressures to be equal to satisfy mechanical equilibrium. For these simulations, we use $N=11,200$ particles in an elongated simulation box oriented along the $z$-axis, and we create enough space in the simulation box to account for both phases, corresponding to a global packing fraction of $\eta^{\mathrm{global}}=0.5184$. The initial particle configuration is arranged in a FCC lattice oriented with the square face perpendicular to the interface, with a crystal packing fraction in the range $\eta^{X}\in[0.5475,0.5575]$. Brownian dynamics simulations are performed for varying packing fractions, incremented by $\Delta\eta^{X}=10^{-4}$. Since the global packing fraction of the system is lower than the initial crystal fraction, the system undergoes phase separation. We simulate this phase separation for $5\times 10^{7}\tau_{\mathrm{BD}}$ using the same time step as that employed for the solid branch of the equation of state. This equilibration period ensures the stabilization of the interfaces between the two coexisting phases. Additionally, we compute the liquid’s equation of state within the range $\eta\in[0.49,0.5]$ in steps of $\Delta\eta=10^{-4}$. We then use the pressure tensor along the $z$-axis, $P_{zz}$, obtained from the coexistence simulations and the equation of state of the liquid $\beta P/\rho$ to calculate the coexistence pressure. The coexistence pressure is obtained by determining the point at which the pressures of the fluid and the fluid-crystal coexistence system are equal. To achieve this, we fit a straight line to the fluid’s equation of state and a second-order polynomial to the pressure tensor results from the coexistence simulations, then apply a root-finding algorithm to solve $\beta P_{zz}/\rho^{X}-\beta P/\rho=0$. To estimate the coexistence pressure and its uncertainty, we employ a bootstrapping method. Specifically, we randomly resample the data with replacement used for fitting both the linear and polynomial curves, generating $n_{b}=10,000$ bootstrap samples. From these samples, we compute the mean and standard deviation, which are reported as the coexistence pressure and its associated uncertainty.
@@ III.1.2 Four dimensional simulations
To perform BD simulations of the four dimensional pHS model, we use an in-house BD code that implements the Ermak-McCammon algorithm [16] extended to four dimensions to solve the BD equations of motion. For the fluid branch of the equation of state, we simulate $N=10000$ particles in the $NVT$ ensemble with a fixed reduced temperature of $k_{B}T/\epsilon=1.4803$ [19]. Using a time step of $\Delta\tau_{\mathrm{BD}}=10^{-5}$, we equilibrate the system for at least $10^{5}\tau_{\mathrm{BD}}$, and we collect data for the pressure during an additional simulation time of $10^{6}\tau_{\mathrm{BD}}$. For the solid branch of the equation of state, we initialize a $D_{4}$ lattice [28, 36] comprised of $N=2048$ particles at the target packing fraction. We equilibrate the system during a simulation time of $10^{6}\tau_{\mathrm{BD}}$, and we collect data during an additional $5\times 10^{6}\tau_{\mathrm{BD}}$; we use the same timestep as the one used for the fluid branch.
For all simulations in either three or four dimensions, standard periodic boundary conditions were used in all directions of the simulation box. However, it is important to note that standard periodic boundary conditions are not efficient or ideal in dimensions $d>3$ since the common cubic lattice packing is not the densest in higher dimensions [37]. Furthermore, non-cubic or non-hypercubic simulation boxes can lead to an efficiency improvement for the computation of nearest neighbors, which helps define neighbor lists for faster neighbor interaction computations.
III.2 Integral equations theory

### v2

@@ Abstract
The hard-sphere potential has become a cornerstone in the study of both molecular and complex fluids. Despite its mathematical simplicity, its implementation in fixed time-step molecular simulations remains a formidable challenge due to the discontinuity at contact. To avoid the issues associated with the ill-defined force at contact, a continuous potential has recently been proposed - here referred to as the pseudo-hard-sphere potential (pHS) [J. Chem, Phys. 149, 164907 (2018)]. This potential is constructed to match the second virial coefficient of the hard-sphere potential and is expected to mimic its thermodynamic properties. However, this hypothesis has only been partially validated within the fluid region of the phase diagram for hard-sphere dispersions in two and three dimensions. In this contribution, we examine the ability of the continuous pHS potential to reproduce the equation of state of a hard-sphere fluid, not only in the fluid phase but also across the fluid-solid coexistence region. Our focus is primarily on the phase diagram of hard-sphere systems in three and four dimensions, however, we also report on the feasibility of the pHS to reproduce the long time dynamics of a three-dimensional colloidal dispersions. We compare the thermodynamic properties obtained from Brownian dynamics simulations of the pHS potential with those derived from refined event-driven simulations of the corresponding hard-sphere potential. Furthermore, we provide a comparative analysis with theoretical equations of state based on both mean-field and integral equation approximations.
@@ I Introduction
The hard-sphere (HS) model remains the reference system for liquids and soft matter due to its simplicity, physical relevance, and rich behavior. It has been extensively studied over the years [1, 2, 3] through computer simulations [4, 5] and is commonly employed as a reference system in perturbation-based thermodynamic approximations [6, 7]. The continued interest in the HS model stems from the extensive knowledge available and the feasibility of experimental realizations [8, 9, 10, 11], made possible by recent advances in microscopy and scattering techniques. These developments, together with complementary computer simulations [12] and theoretical studies using integral equation theory [13, 14], establish the HS model as a robust framework for testing theories across the full spectrum of current soft matter physics research.
@@
Computer simulations are among the principal techniques in soft matter research and have been extensively employed to study the phase diagram of the HS model [15, 12, 3]. However, the inherent discontinuity of the HS model renders continuous time integration methods unsuitable [16, 5]. In particular, Brownian dynamics simulations [17, 2] cannot be directly applied to the HS model, necessitating the use of specialized event-driven algorithms [18]. A mapping of the HS model to a continuous interaction potential was proposed by Jover et al. [19], where the compressibility factor was compared against the well-established Carnahan-Starling (CS) equation of state [2], and the parameters of the potential were fixed when the compressibility factor from simulations matched the analytical CS expression. Rather than relying on trial and error, the work of Báez et al. [20] employed the extended law of corresponding states [21] to determine a set of potential parameters by directly matching the second virial coefficient of the HS potential with that of the continuous interaction potential. This procedure yields a pseudo hard-sphere (pHS) interaction potential with parameters that are independent of thermodynamic conditions, such as density.
@@
The physical criterion proposed by Báez et al. [20] points toward a deeper understanding of the role of the second virial coefficient in explaining the global and local properties of both molecular liquids and soft materials, an aspect that cannot be conceived a priori and therefore it is important to systematically test it. So far, the mapping of the HS model to a continuous potential has been shown to reproduce the thermodynamics of the real HS one-component fluid [19], as well as those of binary and polydisperse mixtures [22]. The pHS model has recently been employed to inverse design the self-assembly of exotic crystalline structures [23]. In addition, computer simulations have been used to investigate the dynamics and transport phenomena of the pHS model [24, 25], and the interaction potential has found applications in modeling active matter [26] and in evaluating depletion forces in binary and ternary colloidal mixtures [27, 28]. These examples underscore the robustness of the model and its wide range of applications in systems where short-ranged, hard-like interactions are required. Furthermore, the pHS interaction potential has been successfully applied within the Ornstein-Zernike framework to reproduce thermodynamic properties, particularly in studies of polydisperse fluids [22]. All these cases make it obvious that, similar to quantum mechanics, there is a group of "isospectral" potentials that lead to identical physical properties. This information is useful not only for proposing a soft and continuous potential that can be used in computer simulation techniques, it highlights the fact that we should explore in more detail the physics that can be extracted from those many-body systems that have the same value of the second virial coefficient.
@@
The aim of this work is to demonstrate the applicability of the pHS model across the full range of the three-dimensional fluid and solid phases and to determine the co-existence densities at the fluid-solid transition using Brownian dynamics simulations. We also provide a brief discussion on the suitability of the pHS potential to replicate the long time dynamics of three-dimensional colloidal dispersions. We solve the Ornstein-Zernike equation with various closure relations to obtain the liquid equation of state over a range of densities. Furthermore, we test the hypothesis proposed by Báez et al. [20] that the potential is applicable in arbitrary spatial dimensions by performing computer simulations and solving the Ornstein-Zernike in four dimensions. Our results show excellent agreement with mean-field equations of state, previously reported simulation data used for fitting semi-phenomenological equations of state, and solutions of the Ornstein-Zernike equation.
@@
Following this Introduction, the manuscript is organized as follows. In Sec. II, we introduce the pHS model, outline its derivation, and specify the parameters used for the interaction potential throughout this work. Sec. III provides a detailed description of the methods employed for the computer simulations and the numerical solution of the Ornstein-Zernike equation, including the relevant parameters. Sec. IV is dedicated to the presentation and discussion of the results obtained from the Ornstein-Zernike equation, focusing mainly on the fluid region. The results of the long-time self-diffusion coefficient as a function of the packing fraction of a three-dimensional colloidal dispersion are discussed in Sec. V. In Sec. VI, we show and examine the computer simulation results for the three- and four-dimensional systems. Finally, Sec. VII concludes the work with an overview of the main findings and perspectives based on the results presented.
@@ II Pseudo-hard sphere potential revisited
The work of Jover et al. [19] introduced a continuous interaction potential, which is a generalized cut-and-shifted Mie potential, to reproduce the CS equation of state. The pHS interaction potential is defined as [19]
$u_{pHS}(r)=\left\{\begin{array}[]{cc}A\varepsilon\left[\left(\frac{\sigma}{r}\right)^{\lambda}-\left(\frac{\sigma}{r}\right)^{\lambda-1}\right]+\varepsilon&r<\sigma B,\\ 0&r\geq\sigma B,\end{array}\right.$ (2)
@@
$A=\lambda\left(\frac{\lambda}{\lambda-1}\right)^{\lambda-1}\quad,\quad B=\frac{\lambda}{\lambda-1}$ (3)
The exponent $\lambda$ is related to the stiffness of the interaction potential, and $\epsilon$ is the energy parameter that measures the repulsion strength between a pair of particles. Typically, $\epsilon$ is used to define the reduced temperature $k_{B}T/\epsilon$, where $k_{B}$ is the Boltzmann constant and $T$ the absolute temperature.
Later, in the work of Báez et al. [20], it was shown that following the so-called extended law of corresponding states [21], the assumption that the second virial coefficient of the continuous potential $B_{2}^{\mathrm{pHS}}$ must be equal to the second virial coefficient of the HS interaction potential, $B_{2}^{\mathrm{HS}}$, will yield a set of values for $\epsilon/k_{B}T$ in terms of $\lambda$ that successfully maps the HS model to a continuous one in any spatial dimension, $d$. In principle, one can choose an arbitrary value for $\lambda$, however, we follow the work of Báez et al. [20] and set the value of $\lambda=50$. For each spatial dimension, a value of the reduced temperature will be obtained. In particular, for three dimensions we use $k_{B}T_{\mathrm{3D}}/\epsilon=1.4737$, and for four dimensions we use $k_{B}T_{\mathrm{4D}}/\epsilon=1.4803$  [20].
@@
In general, the compressibility factor in $d$ dimensions can be written as follows [29],
[…]
where $d\mathbf{r}$ stands for the $d$-dimensional differential volume element. This expression is generic for any interaction potential $u(r)$. However, for the HS interaction potential, the compressibility factor in $d$ dimensions is related to the contact value of the pair correlation function $g(\sigma^{+})$ by [29]
$Z=\frac{\beta P}{\rho}=1+B_{2}\rho g(\sigma^{+})\,,$ (6)
@@
III Computer simulation and integral equation theory details
@@ III.1.1 Three dimensional simulations
For the three dimensional system of particles interacting with the HS potential, we simulate a system of $N=20376$ particles in a fixed volume $V$, and energy $E$. We perform event-driven molecular dynamics (EDMD) using the algorithm of Smallenburg [31]. Initial configurations are obtained by starting in a dilute state at the desired density, and then performing an EDMD simulation in which the particle diameters grow until the desired packing fraction is reached. After the packing fraction was reached, the system is equilibrated for at least $10^{7}\tau_{\mathrm{MD}}$ and data collection for the pressure was done for another $10^{8}\tau_{\mathrm{MD}}$, with $\tau_{\mathrm{MD}}=\sigma\sqrt{m/k_{B}T}$ the simulation time unit, $m$ the mass of the particle, which is the same for all particles. Pressure is measured using the virial expression [31]. For the solid branch, all details remain the same except for the initial configuration, for which a face centered cubic (FCC) lattice was used for each of the target packing fractions. We estimate the average value of the pressure and the error of the measurement using block analysis [32].
@@
The three dimensional pHS model was simulated using the HOOMD-blue code [33]. Brownian dynamics (BD) simulations were performed using the standard Euler scheme to solve the equations of motion for the particles in the canonical ($NVT$) ensemble [17, 34]. For the fluid branch of the equation of state, we simulate a total of $N=32000$ particles at a fixed reduced temperature of $k_{B}T/\epsilon=1.4737$ following the result from Báez et al. [20]. We use a time step of $\Delta t=10^{-5}\tau_{\mathrm{BD}}$, where $\tau_{\mathrm{BD}}=\sigma^{2}/D_{0}$ is the Brownian time unit, and $D_{0}=k_{B}T/3\pi\eta_{0}\sigma$ is the free-particle diffusion coefficient, with $\eta_{0}$ being the zero-frequency shear viscosity. We initialize the system at a low density and slowly compress the simulation box over $10^{5}$ time steps until the target packing fraction has been reached. Once the desired density has been reached, the system is equilibrated for at least $5\times 10^{7}$ time steps, and an additional production run of $10^{7}$ time steps is used to collect data for the pressure, which is computed through the virial expression for a pairwise interaction potential [5]. For the solid branch of the equation of state of the pHS model, we use a smaller time step of $\Delta t=5\times 10^{-7}\tau_{\mathrm{BD}}$, and initialize the system in the FCC configuration for the desired target packing fraction.
[…]
@@
To determine the fluid-crystal coexistence properties in the three dimensional pHS model, we calculate the coexistence pressure using the $NVT$ ensemble method introduced by Smallenburg et al. [35]. Below, we provide a brief summary of the method, but we direct readers to the original publication for a complete description of the method. The coexistence pressure in a fluid-crystal system is estimated by performing direct coexistence simulations in the $NVT$ ensemble to evaluate the system’s pressure tensor. The pressures of the fluid and solid are then compared, as true coexistence requires the pressures to be equal to satisfy mechanical equilibrium. For these simulations, we use $N=11,200$ particles in an elongated simulation box oriented along the $z$-axis, and we create enough space in the simulation box to account for both phases, corresponding to a global packing fraction of $\eta^{\mathrm{global}}=0.5184$. The initial particle configuration is arranged in a FCC lattice oriented with the square face perpendicular to the interface, with a crystal packing fraction in the range $\eta^{X}\in[0.5475,0.5575]$. Brownian dynamics simulations are performed for varying packing fractions, incremented by $\Delta\eta^{X}=10^{-4}$. Since the global packing fraction of the system is lower than the initial crystal fraction, the system undergoes phase separation. We simulate this phase separation for $5\times 10^{7}$ time steps using the same time step as that employed for the solid branch of the equation of state. This equilibration period ensures the stabilization of the interfaces between the two coexisting phases. Additionally, we compute the liquid’s equation of state within the range $\eta\in[0.49,0.5]$ in steps of $\Delta\eta=10^{-4}$. We then use the pressure tensor along the $z$-axis, $P_{zz}$, obtained from the coexistence simulations and the equation of state of the liquid $\beta P/\rho$ to calculate the coexistence pressure. The coexistence pressure is obtained by determining the point at which the pressures of the fluid and the fluid-crystal coexistence system are equal. To achieve this, we fit a straight line to the fluid’s equation of state and a second-order polynomial to the pressure tensor results from the coexistence simulations, then apply a root-finding algorithm to solve $\beta P_{zz}/\rho^{X}-\beta P/\rho=0$. To estimate the coexistence pressure and its uncertainty, we employ a bootstrapping method. Specifically, we randomly resample the data with replacement used for fitting both the linear and polynomial curves, generating $n_{b}=10,000$ bootstrap samples. From these samples, we compute the mean and standard deviation, which are reported as the coexistence pressure and its associated uncertainty.
@@ III.1.2 Four dimensional simulations
To perform BD simulations of the four dimensional pHS model, we use an in-house BD code that implements the Ermak-McCammon algorithm [17] extended to four dimensions to solve the BD equations of motion. For the fluid branch of the equation of state, we simulate $N=10000$ particles in the $NVT$ ensemble with a fixed reduced temperature of $k_{B}T/\epsilon=1.4803$ [20]. Using a time step of $\Delta t=10^{-5}\tau_{\mathrm{BD}}$, we equilibrate the system for at least $10^{5}$ time steps, and we collect data for the pressure during an additional simulation time of $10^{6}$ time steps. For the solid branch of the equation of state, we initialize a $D_{4}$ lattice [29, 37] comprised of $N=2048$ particles at the target packing fraction. We equilibrate the system for $10^{6}$ time steps, and we collect data during an additional $5\times 10^{6}$ time steps; we use the same timestep as the one used for the fluid branch.
For all simulations in three or four dimensions, standard periodic boundary conditions were used in all directions of the simulation box. However, it is important to note that standard cubic periodic boundary conditions are not efficient or ideal in dimensions $d>3$ The conventional cubic periodic boundary condition, which corresponds to a simple cubic tiling, is clearly not optimal, because this sphere packing lattice is not the densest in any  [38] $d\geq 2$. Furthermore, non-cubic or non-hypercubic simulation boxes can lead to an efficiency improvement for the computation of nearest neighbors, which helps to define neighbor lists for faster neighbor interaction computations.
III.2 Integral equation theory